if vectorA=i+j-k and vectorB=2i+3j+pk are mutually perpendicular vectors,find the value of p

Grade:11

2 Answers

gaurav kumar
11 years ago

first find the dot product of the two vectors

A.B = (i+j-k).(2i+3j+pk) = 2+3-p = 0 (Since the dot productof two perpendicular vectors is zero).

 

G Ashwin -
11 years ago

The scalar product to two vectors which are perpendicular is 0.

so,

A.B=2+3-p

A.B=5-p

since scalar product is 0

therefore

5-p=0

p=5